There are 4 charges at the corners of a square. q1 (top left corner)= q. q2 (top right corner)= Q. q3 (bottom left corner) = -q. q4= -Q. Take q = 1.02 µC and Q = 2.23 µC.) The side length of the square is 10 cm.
a) what are the magnitude and direction of the resultant force on q1? (Take q1 as the origin of the coordinate system and measure the angle counterclockwise from the positive x-axis, which is directed towards the right.)
b) What is the resultant force on the center of mass of the four charges?
Answer #1

Data Given

q = 1.02 uC, Q = 2.23 uC, L = 10 cm = 0.10 m

Refer to the diagram

Fox 10 cm FB Fc FNet FD 14.14 cm 10 cm -q -Q 10 cm

F_{B} = \frac{kQq}{BA^{2}}= \frac{9\times 10^{9}\times 1.02\times 10^{-6}\times 2.23\times 10^{-6}}{(0.10)^{2}}= 2.047 N  towards - X axis

F_{D} = \frac{kqq}{AD^{2}}= \frac{9\times 10^{9}\times 1.02\times 10^{-6}\times 1.02\times 10^{-6}}{(0.10)^{2}}= 0.936 N  towards -Y axis

F_{C} = \frac{kqQ}{AD^{2}}= \frac{9\times 10^{9}\times 1.02\times 10^{-6}\times 2.23\times 10^{-6}}{(0.1414)^{2}}= 1.024 N  Along the diameter AC

Now Net Force along X axis will be

F_{x}= F_{Cx}-F_{B} = F_{C}\sin 45^{0}-F_{B}

F_{x}= 1.024 N\sin 45^{0}-2.047 N = -1.323 N

Now Net Force along Y Axis

F_{y}= -F_{Cy}-F_{D} = -F_{C}\cos 45^{0}-F_{D}

F_{y}= -1.024 N\sin 45^{0}-0.936 N = -1.66 N

F_{x}= 1.024 N\sin 45^{0}-2.047 N = -1.323 N

Net Force at A will be

F_{net} = \sqrt{F_{x}^{2}+F_{y}^{2}}= \sqrt{(-1.323)^{2}+(-1.66)^{2}}= 2.123N

Direction will be

\theta = 180^{0}+\tan^{-1}\frac{F_{y}}{F_{x}}= 180^{0}+\tan^{-1}\left ( \frac{-1.66}{-1.323} \right )

\theta = 180^{0}+51.45^{0} = 231.45^{0}  Counterclockwise to + X axis.

Part B) There will be no force at the centre of mass of the four charges as there is no charge is placed.